I have three different problems that I need help on.
The functions f (x) and g(x) = 3x - 2 and g (x) = 2x2. Compute the difference.

If I=PRT and A=P+I, then A =_________.
In two dimensions, the determinant of an ordered group of numbers a, b, c, and d is defined as: ad - bc. Compute the determinant when a, b, c, and d equal 3, -2, 9 and 4 respectively.

Answer by edjones(8007)

1. question is not clear to me.
.
2.
I=prt
p=I/rt
A=p+I
=I/rt + I
=(I+Irt)/rt
.
3.
ad=bc
3*4-(-2*9)
=12-(-18)
=12+18
=30
.
Ed